A car wheel has a mass of 30kg and a radius of 45.0 cm. This happens to be a "spinner" which a thin, solid disk of mass 10kg and a radius of 30 cm on the inside of the wheel. If the car was travelling at 9 m/s, what will be the angular velocity of the "spinner" in the center of the wheel when the car stops? (HINT: assume the car wheel is a hollow cylinder)
Answer #1

initial angular velocity w0 = v/r = 9/0.45 = 20 rad/s

By conservation of angular momentum,

i1w = i1w0 + i2w0 where i1 is moment of inertia of spinner and i2 is that of wheel and w is final angular velocity of spinner

0.5*10*0.30^2*w = 0.5*10*0.30^2*20 + 30*0.45^2*20

w = 290 rad/s answer